OpenStudy (anonymous):

first day attendance is 425 people, day 3 attendance is 575, if attendane is linear increase each day, how many people on day 6

OpenStudy (anonymous):

575-425=150/3=50... 50x6+425=x...725

OpenStudy (anonymous):

I did that wrong...75x6+425=x...875

OpenStudy (anonymous):

Where did 75 come from. I made the same mistake you did, initially!

OpenStudy (anonymous):

you counted day one as part of it, thus it being 150/3 instead of 150/2